OpenStudy (anonymous):

***PLEASE HELP ASAPPPPPPPP, I WILL BECOME A FAN AND GIVE YOU A MEDAL*** This is from Algebra 2: Brad can make 4 key chains in an hour. Velma can make only 3 key chains in an hour, but she already has 6 completed key chains. Explain to Brad how he can use a system of equations to determine when he will have the same number of key chains as Velma.

OpenStudy (greencat):

What's the hurry?

OpenStudy (anonymous):

Due in 3 minutes but I can't figure out what to say.

OpenStudy (greencat):

Is this a test?

OpenStudy (anonymous):

No, worksheet

OpenStudy (greencat):

hang on a sec. gtg for a moment

OpenStudy (anonymous):

....................

OpenStudy (anonymous):

well first Brad will have to determine the amount he has to chatch up to her them you can explian

OpenStudy (anonymous):

yeah true

OpenStudy (greencat):

make equations B=4x V=3x+6 x is hours

OpenStudy (anonymous):

ok so my teacher gave me 2 extra mins. gotta hurry up

OpenStudy (greencat):

just equate the 2 and solve for x

OpenStudy (anonymous):

so basically i just need to plug all of this into words

OpenStudy (greencat):

4x=3x+6 solve

OpenStudy (anonymous):

subtract 6 on both sides then divide by 3

OpenStudy (greencat):

nonono

OpenStudy (greencat):

subtract 3x from both sides

OpenStudy (anonymous):

yeah lol my bad

OpenStudy (anonymous):

and then ivide

OpenStudy (greencat):

it's faster

OpenStudy (anonymous):

divide*

OpenStudy (anonymous):

so basically what can i say

OpenStudy (greencat):

first subract and we'll see

OpenStudy (anonymous):

ok so x = 6

OpenStudy (greencat):

aaand no need to divide now

OpenStudy (anonymous):

yeah ik

OpenStudy (anonymous):

how can i explain how brad used system of equations

OpenStudy (anonymous):

u there?

OpenStudy (greencat):

The system of equations method is different.

OpenStudy (greencat):

just google it. I'm sure purplemath can help

OpenStudy (anonymous):

here is the other part

OpenStudy (greencat):

You get the same answer.

OpenStudy (anonymous):

Explain to Brad how he can use a system of equations to determine when he will have the same number of key chains as Velma.

OpenStudy (anonymous):

hell no this isnt gonna help lmfao

OpenStudy (greencat):

Write the 2 equations: y=4x y=3x+6

OpenStudy (greencat):

Then see if you shuold subtract or add

OpenStudy (anonymous):

yeah and we get 6

OpenStudy (greencat):

Subtract: y=4x minus y=3x-6

OpenStudy (greencat):

no

OpenStudy (greencat):

y=4x -y=-x-6 and add those two (I made the second one negative)

OpenStudy (anonymous):

yeah

OpenStudy (greencat):

 I did part of it

OpenStudy (greencat):

now you do the rest. it's easy

OpenStudy (anonymous):

dude.................................... y = 6...........................

OpenStudy (anonymous):

x = 6

OpenStudy (anonymous):

y=4x y=3x+6 4x = 3x + 6 -3 -3 x = 6